# SpotEdit: Evaluating Visually-Guided Image Editing Methods

### v1

@@ 1 Introduction
Visually-guided image editing enables precise, localized manipulation by combining a reference image with textual instructions to guide generative models. Compared to text-only editing, this multimodal approach provides greater control, stronger semantic alignment, and higher spatial precision, making it valuable for applications such as consistent keyframe editing in the area of long-form video editing (Zhang et al., 2025a; Huang et al., 2025).
Despite rapid advances in diffusion-based (Wu et al., 2025b; Xiao et al., 2025) and autoregressive (Deng et al., 2025; Sun et al., 2024a) generative methods, rigorous evaluation of visually guided editing remains underexplored. Existing benchmarks (Yang et al., 2023; Li et al., 2023) focus largely on coarse manipulations, simple object replacements, or single-object scenes (see Fig. 4), offering limited insight into the complexities of real-world editing. As a result, current evaluations fail to capture the nuanced challenges of multimodal guidance, hindering fair model comparison and progress.
@@
We evaluate leading open- and closed-source models, including OmniGen2 (Zhang et al., 2024), BAGEL (Deng et al., 2025) and GPT-4o (OpenAI, 2025) on SpotEdit. Results reveal that visually guided editing remains fundamentally challenging: the strongest open-source model achieves only $0.685$ similarity score to ground truth. Moreover, models exhibit complementary strengths and weaknesses, e.g., OmniGen2 adheres closely to visual guidance but disrupts background consistency, while BAGEL preserves context yet struggles with cue interpretation. Strikingly, in hallucination cases, even the proprietary GPT-4o performs poorly, hallucinating object presence and executing incorrect edits despite its strong general image editing capabilities.
@@ 2 Related Work
As a pioneer, Paint by Example(Yang et al., 2023) introduced exemplar-based editing by inpainting image regions from a reference exemplar, focusing on visual similarity and identity preservation. DreamEdit (Li et al., 2023) extended this to subject-driven editing, manipulating a subject’s appearance or context while preserving identity. As shown in Fig. 4, prior benchmarks involve simple scenes, few distractors, and near-identical object poses across reference and edited images. In contrast, SpotEdit targets fine-grained, visually-guided editing in complex scenes with multiple objects, sparse textual descriptions, challenging spatial layouts, and pose variations. Its deterministic design with a ground truth enables objective evaluation and, for the first time, tests hallucination to incomplete or missing visual cues.
@@ 3.1 Data Generation Pipeline
We construct the SpotEdit benchmark through a structured data generation pipeline that ingests keyframes extracted from video sources. Our data comes from two complementary datasets: (1) the StoryStream dataset (Yang et al., 2024), a large-scale, high-resolution synthetic multimodal collection designed for long-form story generation, and (2) real (non-synthetic) videos from NExT-QA  (Xiao et al., 2021), which add diversity and visual realism. While StoryStream provides keyframes directly, NExT-QA requires preprocessing to extract them, details of which are given in App. B. Starting with the key-frames, we then utilize our three-stage data generation process, illustrated in Fig. 1 and described below, to generate our benchmark:
@@
Step 2: Frame Localization. To refine the Location of the target object, i.e., specific frames where the target object appears, we query InternVL3-8B (Zhu et al., 2025). Each frame is individually evaluated by the model to determine the presence or absence of the target object.
@@ 4 Experimental Evaluations
Evaluation Setup. For our baseline models, we include the most recent models that support visually-guided image editing. Specifically, UNO (Wu et al., 2025b) and OmniGen (Xiao et al., 2025) are diffusion-based models; BAGEL (Deng et al., 2025) and Emu2 (Sun et al., 2024a) are autoregressive generative models; and OmniGen2 (Zhang et al., 2024) adopts a hybrid architecture that couples an autoregressive text decoder with a diffusion-based image decoder. To compute semantic similarity, we use DINOv2 (Oquab et al., 2023) (and later CLIP (Radford et al., 2021; Cherti et al., 2023) in Tab. 3 of App. C) to extract image representations, as they have been widely adopted in prior works (Wu et al., 2025b, a). Cosine similarity is then computed between the extracted representations, producing scores in the range $0\leq\text{score}\leq 1$.
@@ 4.1 Standard Evaluation
A closer examination of the qualitative examples in Fig. 2 and Fig. 8 confirms that the models indeed produce hallucinations during the editing process. To better quantify hallucination-induced failures, we employ InternVL3-8B(Zhu et al., 2025) as a binary classifier to assess the presence of the target object in the generated image. As shown in Tab. 2, the failure rates are alarmingly high, with GPT-4o emerging as the most vulnerable model in 2 out of the 4 Failure Rate evaluations.
@@ Appendix A Extended Related Work
Recent advances in image editing have introduced models that take an input image along with a textual edit instruction and produce an edited version of the image (Brooks et al., 2023; Yu et al., 2025; Liu et al., 2025; Zhang et al., 2025b; Lin et al., 2025; Wei et al., 2024). However, to provide more concrete guides for image editing one can provide visual guides as wellas textual guide which lead to a more nuanced task called visually-guided image editing. These visual cues help convey richer, more precise editing intent, especially when consistency across multiple frames is required. For example, when editing keyframes in a video, maintaining temporal and stylistic coherence across frames is crucial—something that textual instructions alone cannot achieve. This motivates the use of visually-guided methods, which incorporate reference visuals to better preserve detail and ensure consistency. As expected the methods mentioned before are not capable of performing these tasks. A few recent models have been proposed to address this more complex task.
@@
(visually-guided) Image editing models typically fall into three categories: diffusion models (Xiao et al., 2025; Wu et al., 2025b), autoregressive models (Sun et al., 2024a, b; Deng et al., 2025), and hybrid models (Zhang et al., 2024; Pan et al., 2025) that combine both approaches. Diffusion models including UNO (Wu et al., 2025b) and OmniGen (Xiao et al., 2025) use iterative denoising to produce high-quality images, with UNO focusing on visually-guided controllability and OmniGen unifying tasks like text-to-image generation and editing. Autoregressive models, including BAGEL (Deng et al., 2025), X‑Prompt (Sun et al., 2024b) and Emu2 (Sun et al., 2024a) generate content sequentially via next-token prediction. Bridging the two paradigms, OmniGen2 (Zhang et al., 2024) adopts a hybrid architecture that couples an autoregressive text decoder with a diffusion-based image decoder, allowing for specialized yet coordinated multimodal generation.
Paint by Example (Yang et al., 2023)
DreamEdit (Li et al., 2023)
@@
Paint by Example(Yang et al., 2023) introduced exemplar-based editing by inpainting image regions from a reference exemplar, focusing on visual similarity and identity preservation. DreamEdit (Li et al., 2023) extended this to subject-driven editing, manipulating a subject’s appearance or context while preserving identity. As shown in Fig. 4, prior benchmarks involve simple scenes, few distractors, and near-identical object poses across reference and edited images.
@@
A parallel line of research (Wu et al., 2025a; Liang et al., 2025; Kim et al., 2025) has focused on image composition tasks, where a new image is synthesized by combining elements from multiple context images. Unlike visually-guided editing, these works do not involve editing a specific input image while preserving its structure and style; instead, they aim to generate a novel composition based on the contextual images alone.
@@ Appendix B SpotEdit Details
We construct the SpotEdit benchmark using a structured data generation pipeline, as illustrated in Fig. 1. The pipeline takes as input keyframes extracted from various video sources. For synthetic datasets such as StoryStream(Yang et al., 2024), keyframes are provided directly. For the ShareGPT-4(Xiao et al., 2021) dataset, we employ CLIP-ViT-H-14 (Radford et al., 2021; Cherti et al., 2023) combined with cosine similarity to measure the similarity between consecutive frames. Specifically, the selection process begins with the first frame and progressively includes subsequent frames that have a similarity score less than 0.8 compared to the most recently selected keyframe.
@@
In the first step of the pipeline, the frame captions are provided to the Llama-3.1-8B-Instructhttps://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, which generates the corresponding edit instructions. The prompt used in this step is shown in Fig. 5. For Steps 2 and 3, the instructions are shown in Fig. 1.
@@
(a) Instruction: Replace the red surface in the second image with a flowery surface in the first image.
@@ C.1 Evaluation metrics
  • Background Fidelity: As a more fine-grained evaluation, we assess the model’s ability to preserve the background of the input image while performing the required edit and generating the edited (or output) image. To conduct this evaluation, we first employ GroundingDINO (Liu et al., 2023) to generate bounding boxes around the source object in the input image and the target object in the output image. We then mask out these bounding boxes and compute the similarity score exclusively on the remaining background regions.
  • Object Fidelity: As a complementary evaluation to Background Fidelity, we assess how well the model preserves the identity and appearance of the target object while following the visual guidance from the reference image and applying the edit to the input image. To this end, we employ GroundingDINO (Liu et al., 2023) to extract the target object from both the reference image and the output image. Once isolated, we apply a similarity metric between the two cut-out objects to evaluate how closely they align.
@@ C.2 Comparison to DreamEdit
As discussed in Section 2 and shown in Fig.4, prior benchmarks typically involve simple scenes, few distractors, and nearly identical object poses across the reference and edited images. To quantitatively assess the increased difficulty of our benchmark compared to the previously proposed DreamEdit benchmark(Li et al., 2023), we evaluate models on DreamEdit’s object replacement task, which consists of 198 samples. We use these samples to generate edited images with baseline models and then evaluate their performance using the metrics Global Score, Background Fidelity, and Object Fidelity.
For image similarity computation, we employ the CLIP-ViT-H-14 model (Radford et al., 2021; Cherti et al., 2023) to extract image representations, given its strong performance across a wide range of semantic similarity tasks (Ghazanfari et al., 2024). Cosine similarity is then applied to the representations, yielding scores in the range $0\leq\text{score}\leq 1$.
@@
Emu2 - 0.639 0.679 0.587 $\downarrow$0.09 0.643 0.575 $\downarrow$0.07
OmniGen - 0.429 0.730 0.497 $\downarrow$0.23 0.592 0.431 $\downarrow$0.16
UNO - 0.575 0.599 0.503 $\downarrow$0.10 0.574 0.528 $\downarrow$0.05
BAGEL - 0.672 0.909 0.828 $\downarrow$0.08 0.559 0.496 $\downarrow$0.06
OmniGen2 - 0.719 0.735 0.562 $\downarrow$0.17 0.679 0.697 $\uparrow$0.02

### v2

@@ 1 Introduction
Visually-guided image editing enables precise, localized manipulation by combining a reference image with textual instructions to guide generative models. Compared to text-only editing, this multimodal approach provides greater control, stronger semantic alignment, and higher spatial precision, making it valuable for applications such as consistent keyframe editing in the area of long-form video editing [28, 7].
Despite rapid advances in diffusion-based [21, 23] and autoregressive [3, 17] generative methods, rigorous evaluation of visually guided editing remains underexplored. Existing benchmarks [24, 9] focus largely on coarse manipulations, simple object replacements, or single-object scenes (see Fig. 4), offering limited insight into the complexities of real-world editing. As a result, current evaluations fail to capture the nuanced challenges of multimodal guidance, hindering fair model comparison and progress.
@@
We evaluate leading open- and closed-source models, including OmniGen2 [27], BAGEL [3] and GPT-4o  [^1] https://openai.com/index/introducing-4o-image-generation/ on SpotEdit. Results reveal that visually guided editing remains fundamentally challenging: the strongest open-source model achieves only $0.685$ similarity score to ground truth. Moreover, models exhibit complementary strengths and weaknesses, e.g., OmniGen2 adheres closely to visual guidance but disrupts background consistency, while BAGEL preserves context yet struggles with cue interpretation. Strikingly, in hallucination cases, even the proprietary GPT-4o performs poorly, hallucinating object presence and executing incorrect edits despite its strong general image editing capabilities.
@@ 2 Related Work
As a pioneer, Paint by Example[24] introduced exemplar-based editing by inpainting image regions from a reference exemplar, focusing on visual similarity and identity preservation. DreamEdit [9] extended this to subject-driven editing, manipulating a subject’s appearance or context while preserving identity. As shown in Fig. 4, prior benchmarks involve simple scenes, few distractors, and near-identical object poses across reference and edited images. In contrast, SpotEdit targets fine-grained, visually-guided editing in complex scenes with multiple objects, sparse textual descriptions, challenging spatial layouts, and pose variations. Its deterministic design with a ground truth enables objective evaluation and, for the first time, tests hallucination to incomplete or missing visual cues.
@@ 3.1 Data Generation Pipeline
We construct the SpotEdit benchmark through a structured data generation pipeline that ingests keyframes extracted from video sources. Our data comes from two complementary datasets: (1) the StoryStream dataset [25], a large-scale, high-resolution synthetic multimodal collection designed for long-form story generation, and (2) real (non-synthetic) videos from NExT-QA  [22], which add diversity and visual realism. While StoryStream provides keyframes directly, NExT-QA requires preprocessing to extract them, details of which are given in App. B. Starting with the key-frames, we then utilize our three-stage data generation process, illustrated in Fig. 1 and described below, to generate our benchmark:
@@
Step 2: Frame Localization. To refine the Location of the target object, i.e., specific frames where the target object appears, we query InternVL3-8B [30]. Each frame is individually evaluated by the model to determine the presence or absence of the target object.
@@ 4 Experimental Evaluations
Evaluation Setup. For our baseline models, we include the most recent models that support visually-guided image editing. Specifically, UNO [21] and OmniGen [23] are diffusion-based models; BAGEL [3] and Emu2 [17] are autoregressive generative models; and OmniGen2 [27] adopts a hybrid architecture that couples an autoregressive text decoder with a diffusion-based image decoder. To compute semantic similarity, we use DINOv2 [14] (and later CLIP [16, 2] in Tab. 3 of App. C) to extract image representations, as they have been widely adopted in prior works [21, 20]. Cosine similarity is then computed between the extracted representations, producing scores in the range $0\leq\text{score}\leq 1$.
@@ 4.1 Standard Evaluation
A closer examination of the qualitative examples in Fig. 2 and Fig. 8 confirms that the models indeed produce hallucinations during the editing process. To better quantify hallucination-induced failures, we employ InternVL3-8B[30] as a binary classifier to assess the presence of the target object in the generated image. As shown in Tab. 2, the failure rates are alarmingly high, with GPT-4o emerging as the most vulnerable model in 2 out of the 4 Failure Rate evaluations.
@@ Appendix A Extended Related Work
Recent advances in image editing have introduced models that take an input image along with a textual edit instruction and produce an edited version of the image [1, 26, 13, 29, 11, 19]. However, to provide more concrete guides for image editing one can provide visual guides as wellas textual guide which lead to a more nuanced task called visually-guided image editing. These visual cues help convey richer, more precise editing intent, especially when consistency across multiple frames is required. For example, when editing keyframes in a video, maintaining temporal and stylistic coherence across frames is crucial—something that textual instructions alone cannot achieve. This motivates the use of visually-guided methods, which incorporate reference visuals to better preserve detail and ensure consistency. As expected the methods mentioned before are not capable of performing these tasks. A few recent models have been proposed to address this more complex task.
@@
(visually-guided) Image editing models typically fall into three categories: diffusion models [23, 21], autoregressive models [17, 18, 3], and hybrid models [27, 15] that combine both approaches. Diffusion models including UNO [21] and OmniGen [23] use iterative denoising to produce high-quality images, with UNO focusing on visually-guided controllability and OmniGen unifying tasks like text-to-image generation and editing. Autoregressive models including BAGEL [3], X‑Prompt [18] and Emu2 [17] generate content sequentially via next-token prediction. Bridging the two paradigms, OmniGen2 [27] adopts a hybrid architecture that couples an autoregressive text decoder with a diffusion-based image decoder, allowing for specialized yet coordinated multimodal generation. More recently, a new paradigm has emerged that unifies visual understanding [4, 6] with generation, with BAGEL [3] being a representative example of this category.
Paint by Example [24]
DreamEdit [9]
@@
Paint by Example[24] introduced exemplar-based editing by inpainting image regions from a reference exemplar, focusing on visual similarity and identity preservation. DreamEdit [9] extended this to subject-driven editing, manipulating a subject’s appearance or context while preserving identity. As shown in Fig. 4, prior benchmarks involve simple scenes, few distractors, and near-identical object poses across reference and edited images.
@@
A parallel line of research [20, 10, 8] has focused on image composition tasks, where a new image is synthesized by combining elements from multiple context images. Unlike visually-guided editing, these works do not involve editing a specific input image while preserving its structure and style; instead, they aim to generate a novel composition based on the contextual images alone.
@@ Appendix B SpotEdit Details
We construct the SpotEdit benchmark using a structured data generation pipeline, as illustrated in Fig. 1. The pipeline takes as input keyframes extracted from various video sources. For synthetic datasets such as StoryStream[25], keyframes are provided directly. For the ShareGPT-4[22] dataset, we employ CLIP-ViT-H-14 [16, 2] combined with cosine similarity to measure the similarity between consecutive frames. Specifically, the selection process begins with the first frame and progressively includes subsequent frames that have a similarity score less than 0.8 compared to the most recently selected keyframe.
@@
In the first step of the pipeline, the frame captions are provided to the Llama-3.1-8B-Instruct [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, which generates the corresponding edit instructions. The prompt used in this step is shown in Fig. 5. For Steps 2 and 3, the instructions are shown in Fig. 1.
@@
(c) Instruction: Match the writing on the man’s shirt in the second image with the first image.
@@ C.1 Evaluation metrics
  • Background Fidelity: As a more fine-grained evaluation, we assess the model’s ability to preserve the background of the input image while performing the required edit and generating the edited (or output) image. To conduct this evaluation, we first employ GroundingDINO [12] to generate bounding boxes around the source object in the input image and the target object in the output image. We then mask out these bounding boxes and compute the similarity score exclusively on the remaining background regions.
  • Object Fidelity: As a complementary evaluation to Background Fidelity, we assess how well the model preserves the identity and appearance of the target object while following the visual guidance from the reference image and applying the edit to the input image. To this end, we employ GroundingDINO [12] to extract the target object from both the reference image and the output image. Once isolated, we apply a similarity metric between the two cut-out objects to evaluate how closely they align.
@@ C.2 Comparison to DreamEdit
As discussed in Section 2 and shown in Fig.4, prior benchmarks typically involve simple scenes, few distractors, and nearly identical object poses across the reference and edited images. To quantitatively assess the increased difficulty of our benchmark compared to the previously proposed DreamEdit benchmark[9], we evaluate models on DreamEdit’s object replacement task, which consists of 198 samples. We use these samples to generate edited images with baseline models and then evaluate their performance using the metrics Global Score, Background Fidelity, and Object Fidelity.
For image similarity computation, we employ the CLIP-ViT-H-14 model [16, 2] to extract image representations, given its strong performance across a wide range of semantic similarity tasks [5]. Cosine similarity is then applied to the representations, yielding scores in the range $0\leq\text{score}\leq 1$.
@@
[…]